OpenStudy (anonymous):

WILL GIVE MEDAL,Which mixed number is equal to 80.78?

ganeshie8 (ganeshie8):

As a start, split the given number into integer and fractional parts : \[80.78 = 80 + 0.78 \]

ganeshie8 (ganeshie8):

next work on the fractional part try multiplying it by \(\frac{100}{100}\)

OpenStudy (anonymous):

how

ganeshie8 (ganeshie8):

\[\begin{align}80.78 &= 80 + 0.78 \\~\\ &= 80 + 0.78 \times \dfrac{100}{100} \\~\\ &= 80 + \dfrac{78}{100} \end{align}\]

ganeshie8 (ganeshie8):

cancel any common factors in that fraction ^

ganeshie8 (ganeshie8):

looks 2 goes in both 78 and 100, eh ?

ganeshie8 (ganeshie8):

\[\require{cancel}\begin{align}80.78 &= 80 + 0.78 \\~\\ &= 80 + 0.78 \times \dfrac{100}{100} \\~\\ &= 80 + \dfrac{78}{100}\\~\\ &= 80 + \dfrac{\cancel{78}^{39}}{\cancel{100}^{50}}\\~\\ &= 80 + \dfrac{39}{50}\\~\\ &= 80 \dfrac{39}{50}\\~\\ \end{align}\]

ganeshie8 (ganeshie8):

see if that makes more or less sense